OpenStudy (anonymous):

how to work out 30th term of 8,11,14,17?

OpenStudy (jamesj):

What's the fifth term?

OpenStudy (anonymous):

20

OpenStudy (jamesj):

So how did you get that?

OpenStudy (anonymous):

add 3

OpenStudy (jamesj):

exactly. So if the first term x1 = 8, then x2 = 11 = x1 + 3 x3 = x1 + 3 + 3 = x1 + (3-1)*3 x4 = x1 + (4-1)*3 so, what's the expression for the 30th term?

OpenStudy (anonymous):

t30 = 1 + (30-1)3 = 88 but when i do it manually by adding 3 it comes up as 95 how do i work it out so it comes to 95

OpenStudy (jamesj):

x30 = x1 + (30-1)*3, where x1 = 8

OpenStudy (anonymous):

this series forms an Arithmetic PRogression... where first term is 8 and common difference is 30 T30 = 8 + (30-1)*3 T30 = 8 + 87 T30 = 95.. is this correct ??

OpenStudy (anonymous):

yh thanks man

OpenStudy (anonymous):

x = a+ (n-1)d x=8+(30-1)3 8+87=95

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

what is first 6 terms in the expansion of 1/(1-3x)^5

OpenStudy (anonymous):

can someone help me on this i dont get it

OpenStudy (amistre64):

what are the values when x=(1,2,3,4,5 and 6)